Bags of chips are filled using a machine. Because not al chips are the same size, it is impossible to get the exact target weight in each bag. The target weight is 65 g. Based on samples of size 4, the average standard deviation is about 0.5 g, and the average weight of the last batch of four bags was 64 g. Which of the following are the upper and lower control limits for an x control chart? (62.5, 65.5) 9 (64.25, 64.75) 9 (63.5, 66.5) 9 o (64.25, 65.75) g
Answer #1

here upper control limit =target weight +3*std deviation/sqrt(n)=65+3*0.5/sqrt(4)=65.75

and lower control limit =target weight -3*std deviation/sqrt(n)=65-3*0.5/sqrt(4)=64.25

therefore correct option is : (64.25; 65.75 ) g